In a survey conducted by the American Animal Hospital Association, 37 % of the respondents stated that they talk to their pets on the phone. Pagotes is a veterinarian and he doubts this. He thinks it’s less. He randomly selects 150 pet owners and discovered that 54 of them spoke to their pet on the telephone. Can Pagotes be sceptical? Analyse the problem at the 5% level of significance.
Answer #1

Given P. - . 37 n= iso x = 54 P = P. Us HA :P <Po 54 P- 2 h 0.36 Iso As 0.37 2. = 1 - 0.37 - 0263 test statistic = 2 = P-P. u